Q3. To make a beef soup base you want to boil 1kg of beef ( assume pure protein) in 5L of water starting at STP (a)How much heat in Joules do you need to apply? (b) You left it boiling and forgot it, all the water boiled out to steam at 100 °C. Ignoring the beef in the calculations, how much additional heat in Joules was used In your calculations ignore the container in which you boil
Answer #1

(a) as 1 ltr = 10-3m3

So total water volume = 5 × 10-3 m 3

  Hence mass m = volume × density = 5 × 10-3 × 103 = 5 kg

Specific heat capacity of water c = 4200 J/kg

Temperature at STP = 230C and boiling temperature = 100

Hence ∆T = 100 - 23 = 770C

Hence heat required Q = mc∆T = 5 × 4200 × 77 = 1617 kJ

(b) to vaporize the water Q = mL

Where L is the latent heat of vaporization

Hence Q = 5 × 540 × 4200 = 11340 kJ

Hence additional heat = 11340 - 1617 = 9723 kJ